A mass of 165 kg is hanging on an extension of the pole(L-shaped pole) and i want to calculate the forces of distribution on various points of the pole so that it will not collapse
1
Expert's answer
2019-12-25T12:14:14-0500


Static law


"R-F=0"

Torque law


"M-F\cdot l=0"

"F=165\cdot 10=1650 (N)"